Consider the equilibrium reaction: AX3(s) ⇌ A3+(aq) + 3X-(aq). What is the solubility of AX₃ (MM = 78.0 g/mol) in mg/L at 25⁰C if its Ksp is 1.0 x 10⁻³³?
1
Expert's answer
2021-01-04T08:03:40-0500

"AX_{3(s)} \u21cc A^{3+}_{(aq)} + 3X^-_{(aq)}"

Ksp = 1.0 x 10⁻³³


Let us denote the solubility of AX3 as S molL–1.

[A3+] = S

[X-] = 3S


Ksp = [A3+][X-]³ = [S][3S]³ = 27S⁴

1.0 x 10⁻³³ = 27S⁴

S = 2.47×10-⁹ M


Molar mass of AX3 = 78.0 g/mol

S = 2.47 × 10-9 moldm-3 × 78.0 g/mol = 1.93 × 10-7 gdm-3 = 0.193 mg per 1L


Therefore, the the solubility of AX₃ at 25⁰C is 0.193 mg/L